Students in a school agreed to buy a trophy that would be presented to the winning soccer team. Before the trophy was purchased, 10 students left the school, and this meant that all other students had to pay an extra dollar. It was then decided that 15 students who did not play soccer should not contribute to the cost of the trophy. This resulted in the remaining students' cost increasing another $2 each. If the trophy cost $n, in which range is n?
Answer by richwmiller(17219)

s*c=n
(s-10)(c+1)=n
(s-25)(c+3)=n
three equations three unknowns